OpenStudy (anonymous):

i'm not sure where to start i think its something like this 13x=34 15x=x>34 is that right? A company manufactures a product that has a unit selling price of $15 and a unit cost of $13. If fixed costs are $34, determine the number of units that must be sold in order for the company to have a profit.

OpenStudy (amistre64):

profit = money in - money out

OpenStudy (amistre64):

also none as revenue - costs

OpenStudy (amistre64):

revenue = number sold * price per item

OpenStudy (amistre64):

and you want the "break even price right?

OpenStudy (amistre64):

13x + 34 = 15x should be the break even point

OpenStudy (amistre64):

you sell: 13(1) + 34 ?= 15(1) 47 > 15 costs are greater then revenue 13(2) + 34 ?= 15(2) 26 + 34 ?= 30 60 > 30 y = 2x - 34 is the graph; when that is above 0 you make a profit i believe

OpenStudy (amistre64):

13(17) + 34 ?= 15(17) 221 + 34 ?= 255 255 = 255 yep, gotta sell 17 to break even